# A dark-field setup for the measurement of light-by-light scattering with high-intensity lasers

### v1

@@ II.1 Theoretical Basics
$N_{\parallel,\perp}\sim c_{\parallel,\perp}\,\alpha^{4}\bigl{(}1-\cos\vartheta_{\rm coll}\bigr{)}^{4}\Bigl{(}\frac{W_{\rm pump}}{m}\frac{\omega_{\rm probe}}{m}\Bigr{)}^{2}N_{\rm probe}\,,$ (6)
with $\alpha$, the collision angle $\vartheta_{\rm coll}$, the pulse energy of the pump $W_{\rm pump}$, the probe photon energy $\omega_{\rm probe}$ and the probe photon number $N_{\rm probe}\simeq W_{\rm probe}/\omega_{\rm probe}$. Finally, the coefficients $c_{\parallel,\perp}$ governing the strength of the signal components induced in a $\parallel$,$\perp$ , ⟂-polarized mode depend on both the low-energy constants (2) and the polarizations of the driving laser fields. Besides, the signal photon numbers typically depend non-trivially on other characteristic parameters of the colliding beams such as their pulse durations, waist sizes and Rayleigh ranges; cf., e.g., Ref. Mosman:2021vua . In the paraxial limit, this function amounts to a global factor in Eq. (6) in the sense that it is independent of the polarization characteristics of the driving beams as well as the signal. In turn, ratios of the explicit values of $c_{\parallel,\perp}$ provided below reproduce the ratios of the associated signal photon numbers. Note, however, that higher-order corrections in the diffraction angles characterizing both the colliding beams and the signal, which can for instance be safely neglected for state-of-the-art XFEL probes, generically break this factorization. Besides, beyond the low-energy limit also higher-order derivative corrections Gusynin:1995bc ; Gusynin:1998bt ; Karbstein:2021obd to Eq. (1) are expected to impact the signal polarizations King:2023eeo . Because in the present context we are only interested in an estimate for the signals to provide for a reference and guidance for our full numerical calculations detailed below, in our analytical calculations we restrict ourselves to the paraxial limit. In this case, the coefficients $c_{\parallel,\perp}$ in Eq. (6) are given by
@@
A comparison of the results in Eqs. (7)-(9) then implies that the maximum signal photon number is obtained for linearly polarized pump and probe beams and $\phi=\pi/2$ for which $c_{\parallel}+c_{\perp}\simeq 4b^{2}$. Because $c_{\perp}\simeq 0$ in this case, this number also amounts to the maximum possible signal polarized similarly to the probe. On the other hand, the maximum polarization/helicity-flip signal is obtained in the collision of linearly polarized pump and probe beams with $\phi=\pi/4$ or a circularly polarized probe collided with a linearly polarized pump, respectively. Both of these cases result in $c_{\perp}=(a-b)^{2}$.
@@ II.2 Experimental Setup
We address this challenge by generating a double pulse each containing an energy of several tens of Joules. The two pulses will be introduced in the frontend of the high-intensity laser system by splitting the original pulse, introducing an appropriate temporal delay of about $2\,{\rm ns}$. The delay and the pulse energy content can be finely tuned by a delay stage and an appropriate continuous Neutral Density (ND) filter. The two pulses will then be recombined and subsequently travel the exact same path through the amplifiers, compressor and the laser beam delivery to the experimental vacuum chamber, where a $90^{\circ}$ off-axis parabolic mirror (OAP) with effective focal length of $f_{\rm eff}=30\,{\rm cm}$ focuses the pulses. A spherical mirror with suitable radius of curvature ($R=30\,{\rm cm}$) will be mounted downstream of the primary focus point. It will reflect and refocus the pulses, such that the first pulse on its return collides with the second pulse in its first pass in the common focus spot. A motorized hexapod will enable precise spatial overlap, whereas fine tuning the temporal delay between the pulses in the frontend varies temporal overlap.
@@ Appendix A Components of $\vec{E}(\omega,\vec{k})$ and far-field polarization vectors
where the arguments $\tilde{\rm x}$ and $\tilde{\rm y}$ of $E_{0,{\rm x}}$ and $E_{0,{\rm y}}$ are to be identified with Eq. (23). One can easily convince oneself that the electric field components in Eq. (31) indeed fulfill $\vec{k}\cdot\vec{E}(\omega,\vec{k})=0$ in accordance with Maxwell’s equations in vacuum. The unitary vector

### v2

@@ II.1 Theoretical Basics
$N_{\parallel,\perp}\sim c_{\parallel,\perp}\,\alpha^{4}\Bigl{(}\frac{W_{\rm pump}}{m}\frac{\omega_{\rm probe}}{m}\Bigr{)}^{2}N_{\rm probe}\,,$ (6)
with $\alpha$, the pulse energy of the pump $W_{\rm pump}$, the probe photon energy $\omega_{\rm probe}$ and the probe photon number $N_{\rm probe}\simeq W_{\rm probe}/\omega_{\rm probe}$. Finally, the coefficients $c_{\parallel,\perp}$ governing the strength of the signal components induced in a $\parallel$,$\perp$ , ⟂-polarized mode depend on both the low-energy constants (2) and the polarizations of the driving laser fields. Besides, the signal photon numbers typically depend non-trivially on other characteristic parameters of the colliding beams such as the collision angle, their pulse durations, waist sizes and Rayleigh ranges; cf., e.g., Ref. Mosman:2021vua ; Karbstein:2021ldz . Specifically, the dependence on the collision angle is such that the signal vanishes for co-propagating beams and becomes maximum for counter-propagation. In the paraxial limit, these additional dependencies are encoded in a function that amounts to global factor in Eq. (6) in the sense that it is independent of the polarization characteristics of the driving beams as well as the signal. In turn, ratios of the explicit values of $c_{\parallel,\perp}$ provided below reproduce the ratios of the associated signal photon numbers. Note, however, that higher-order corrections in the diffraction angles characterizing both the colliding beams and the signal, which can for instance be safely neglected for state-of-the-art XFEL probes, generically break this factorization. Besides, beyond the low-energy limit also higher-order derivative corrections Gusynin:1995bc ; Gusynin:1998bt ; Karbstein:2021obd to Eq. (1) are expected to impact the signal polarizations Dinu:2013gaa ; King:2023eeo . Because in the present context we are only interested in an estimate for the signals to provide for a reference and guidance for our full numerical calculations detailed below, in our analytical calculations we restrict ourselves to the paraxial limit. In this case, the coefficients $c_{\parallel,\perp}$ in Eq. (6) are given by
@@
A comparison of the results in Eqs. (7)-(9) then implies that the maximum signal photon number is obtained for linearly polarized pump and probe beams and $\phi=\pi/2$ Karbstein:2019bhp for which $c_{\parallel}+c_{\perp}\simeq 4b^{2}$. Because $c_{\perp}\simeq 0$ in this case, this number also amounts to the maximum possible signal polarized similarly to the probe. On the other hand, the maximum polarization/helicity-flip signal is obtained in the collision of linearly polarized pump and probe beams with $\phi=\pi/4$ Aleksandrov:1985 or a circularly polarized probe collided with a linearly polarized pump, respectively. Both of these cases result in $c_{\perp}=(a-b)^{2}$.
@@ II.2 Experimental Setup
We address this challenge by generating a double pulse each containing an energy of several tens of Joules. The two pulses will be introduced in the frontend of the high-intensity laser system by splitting the original pulse, introducing an appropriate temporal delay of about $2\,{\rm ns}$. The delay and the pulse energy content can be finely tuned by a delay stage and an appropriate continuous Neutral Density (ND) filter. Subsequently these pulses will travel the exact same path through the amplifiers, compressor and the laser beam delivery to the experimental vacuum chamber, where a $90^{\circ}$ off-axis parabolic mirror (OAP) with effective focal length of $f_{\rm eff}=30\,{\rm cm}$ focuses the pulses. A spherical mirror with suitable radius of curvature ($R=30\,{\rm cm}$) will be mounted downstream of the primary focus point. It will reflect and refocus the pulses, such that the first pulse on its return collides with the second pulse in its first pass in the common focus spot. A motorized hexapod will enable precise spatial overlap, whereas fine tuning the temporal delay between the pulses in the frontend varies temporal overlap. The spatial overlap between the pulses can be optimized to better than $<0.1\,\upmu{\rm m}$ with typically available hardware. At the same time, fine tuning the temporal delay between the pulses in the frontend to vary the temporal overlap is possible to better than ${\rm fs}$ precision. We note, however, that this overlap will be influenced by shot-to-shot variations, especially of the beam pointing, present in any large laser system. In principle, these can be minimized by active control to a similar or better level. The necessary level of improvement will depend on the real-world background contributions and thus is beyond the scope of this manuscript.
@@ Appendix A Components of $\vec{E}(\omega,\vec{k})$ and far-field polarization vectors
where the arguments $\tilde{\rm x}$ and $\tilde{\rm y}$ of $E_{0,{\rm x}}$ and $E_{0,{\rm y}}$ are to be identified with Eq. (23). One can easily convince oneself that the electric field components in Eq. (31) indeed fulfill $\vec{k}\cdot\vec{E}(\omega,\vec{k})=0$ in accordance with Maxwell’s equations in vacuum. The unit vector